The digits of a three-digit number add up to 18. If the ten’s digit is twice the hundred’s digit and the hundred’s digit is 1/3 the unit’s digit, what is the number? ?

Answer by mananth(16946)

x+y+z=18
y=2x
x=z/3
..
Plug the value of y & z in the first equation
x+2x+3x=18
6x=18
x=3
..
y=2x
y=6
..
z=3x
z=3*3=9
The number is 369